[Brahms-l] Shift report 20040113 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Tue Jan 13 2004 - 15:30:12 EST
------------------------------------------------------------
Date: Tue Jan 13 15:30:11 2004, Shift: 08:00-16:00
Supervisor: Sorin ZGURA, with: Ionut ARSENE
Log book#15 pages 33 - 33

Shift Summary:
------------------------------------------------------------
1. Some problems with injection. Beam in the last part of the shift.
2. 9663 - Non-TPC pedestals
   9664 - TPC pedestals
   9667 - Data run (FS@12 deg. 1/8 field; MRS@90 deg. same field)
   9668 - Data run (same settings)
   9669 - Data run (same settings)

3. Problems with H2; data was taken without it.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Jan 13 15:30:27 2004

This archive was generated by hypermail 2.1.8 : Tue Jan 13 2004 - 15:30:43 EST